OpenStudy (anonymous):

If y varies directly as x, find the constant of variation if y = 32 when x = 8. A. 12 B. 4 C. 256 D. 3

OpenStudy (anonymous):

B

OpenStudy (paki):

If Y varies directly with x that means that the ratio of Y/x= a constant... then the k will be equal to... Y/x=k 8/2=k 4=k @musa19965